Trials in an experiment with a polygraph include 97 results that include 23 cases of wrong results and 74 cases of correct results. Use a 0.01 significance level to test the claim that such polygraph results are correct less than 80​% of the time. Identify the null​ hypothesis, alternative​ hypothesis, test​ statistic, P-value, conclusion about the null​ hypothesis, and final conclusion that addresses the original claim. Use the​ P-value method. Use the normal distribution as an approximation of the binomial distribution.
Answer #1

Solution:

n = 97

x = 74

1 - x = 23

\hat p = x/n = 74/97 = 0.7629

Claim: p < 80% i.e. p < 0.80

The null and alternative hypothesis are

H0 : p = 0.80

H1 : p < 0.80

The test statistic z is

z =   - P(1-

= (0.7629 - 0.80)/\sqrt{}[0.80*(1 - 0.80)/97]

= -0.91

Test statistic z = -0.91

For left tailed test :

p value = P(Z < z)

= P(Z < -0.91)

= 0.1814 (use z table)

p value = 0.1814

Use a 0.01 significance level .

p value is greater than 0.01

Decision : Fail to reject the null hypothesis,

Conclusion: Not sufficient evidence to support the claim that p is less than 80%